–3x + y = 2
–9x + 3y = 6

Answer by checkley79(3341)

–3x + y = 2 MULTIPLY BY -3 & ADD.
–9x + 3y = 6
9X-6Y=-6
---------------
-3Y=0
Y=O ANS.
-3X+0=2
X=2/-3=-2/3 ANS.
PROOF:
-9(-2/3)+3*9=6
18/3+0=6
6=6